OpenStudy (anonymous):

If there is a 5% chance that my Heavy will die, a 10% chance that your Medic will die and if these are independent events, what is the probability that we both stay alive? A) 0.05 or 5% B) 0.5 or 50% C) 0.855 or 85.5% D) 0.9 or 90%

pooja195 (pooja195):

5% = 1/20, 10% = 1/10 (19/20) x (9/10) = 171/200 or 85.5%

OpenStudy (anonymous):

So it would be 5%?

jimthompson5910 (jim_thompson5910):

no, pooja195 has it correct

pooja195 (pooja195):

WELCOME TO OPENSTUDY!

OpenStudy (anonymous):

So it would be 85.5%

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok thanks